A firms income statement included the following data. The firms average tax rate was 20% $ 9,400 Cost of goods sold Income
Answer #1
a] Firm's net income = 3400*80%/20% = $            13,600
b] Firm's revenue = 13600+3400+2400+2400+4400+9400 = $            35,600
c] EBIT = 35600-9400-4400-2400 = $            19,400
CANDY CANES: April May June
Sales $                     -   $       3,30,000
Net income $                     -   $           68,000
Net cash flow $      -2,62,000 $                    -   $      3,30,000